# fftvis: A Non-Uniform Fast Fourier Transform Based Interferometric Visibility SimulatorA

### v1

@@ 5.2 Memory Requirements
    NUFFT Intermediate Grid: The temporary uniform grid used internally by finufft. Its size in each dimension $i$, $n_{i}$, is determined by Equation 5, which depends on the maximum source extent (which we take to be $X_{i}=1$ for sky models that extend to the horizon), the maximum baseline coordinate, $S_{i}$, the desired precision, $\epsilon$, and oversampling factor, $\sigma$. The memory explicitly scales as $4\sigma N_{\rm grid}$.
@@
In matvis, a practical approach to managing memory involves chunking computations along the source axis. This technique allows for a nearly arbitrary reduction in the peak memory usage by processing subsets of sources sequentially, with generally minimal computational overhead. While the overhead might increase for very fine-grained chunking, it typically remains manageable for current hardware memory capacities. However, this source chunking strategy is generally less useful for fftvis when the maximum extent of the array is large. As shown in the memory tests above, the memory usage of fftvis is typically dominated by the size of the intermediate FFT grid. This grid size is not dependent on the number of sources, but rather on the maximum angular extent of the sky model being considered ($X_{i}$) and the maximum baseline coordinates ($S_{i}$). As a result, simply reducing the number of sources processed per chunk does not shrink the required grid size if the sources across all chunks collectively span the same wide angular extent. One potential way to mitigate poor memory scaling in fftvis would involve spatial chunking, or dividing the simulations into chunks based on the sources occupying localized regions of the sky. By limiting the angular extent ($X_{i}$) for sources within a single chunk, the factors determining the grid size in Equation 5 could be reduced, thus decreasing the maximum memory required for processing that chunk. Implementing such a scheme would require careful partitioning of the source catalog and potentially reinitializing the NUFFT for each spatial chunk, which may increase the computational runtime. We leave the development and evaluation of this spatial chunking method for future work.
@@
Figure 7: Comparison of the estimated memory usage of fftvis and matvis for the computational scaling cases presented in Figure 5. For results shown here, we assume no chunking of the sky model in the processing stage, which can reduce the memory usage for matvis by summing over subsets of the sky model. We find that memory usage of fftvis is significantly less than that of matvis especially when the number of antennas and sources is large due to matvis forming an $N_{\rm sources}N_{\rm ants}$ matrix. When $u_{\max}$ is large, fftvis memory usage exceeds that of matvis due to the formation of the underlying fine FFT grid that scales with the physical size of the array in units of wavelengths. As with the computational scaling explored in Figure 5, the memory usage for arrays with a large $u_{\rm max}$ can be dramatically reduced if the array is gridded, as shown at in the right-hand column.
@@ 5.3 Multicore Performance and Parallelization
In fftvis, we adopt a hybrid strategy that prioritizes multiprocessing while still incorporating multithreading. Specifically, we parallelize the entire workflow across multiple processes, then assign the remaining cores as threads within each process. This approach is motivated by findings in Kittiwisit et al. (2025), where a pure multithreaded strategy exhibited a performance plateau after just a few cores due to inherent scaling limitations and shared resource contention. In contrast, a multiprocessing strategy avoids many of these bottlenecks, providing more favorable scaling across a larger number of cores, albeit with a modest overhead associated with initializing processes.
To manage the distribution of these independent tasks across processes on a single machine, fftvis takes advantage of the ray library. Originally developed for scalable machine learning, ray provides a robust framework for efficiently scaling Python applications from a single core to many. Specifically, fftvis uses ray to divide the overall simulation workload by distributing the calculations across the $N_{\rm proc}$ available processes, assigning each process a roughly equal portion of the total workload determined by the number of simulation time steps and frequency channels. This approach helps manage the creation and coordination of multiple processes, simplifying the implementation of our chosen multiprocessing strategy. Furthermore, ray supports shared memory, enabling large, read-only objects like the beam, sky model, and model coordinates to be loaded once and accessed by multiple worker processes. While this shared memory capability is beneficial, a current limitation within fftvis’s implementation is that each process still independently interpolates the beam and performs coordinate rotations for all $N_{\rm sources}$ on each process independently. Therefore, these specific steps do not fully utilize memory sharing, requiring that arrays of size $\mathcal{O}\left(N_{\rm sources}\right)$ be stored in memory for each independent process.
@@ 6 Conclusion
We have demonstrated through a series of validation tests that fftvis achieves numerical accuracy comparable to the established visibility simulator, matvis, itself validated against pyuvsim in Kittiwisit et al. (2025). Our first set of comparisons against analytic solutions for diffuse sky models confirms the accuracy of the NUFFT implementation, while assessments of delay and fringe-rate spectra show that fftvis preserves the spectral and temporal coherence required for end-to-end validation tests. We find that the visibilities produced do not introduce significant spectral or temporal artifacts when operated at sufficient NUFFT precision ($\epsilon\lesssim 10^{-13}$). Furthermore, our benchmarking tests highlight the specific regimes where fftvis offers substantial advantages. For simulations involving large, densely-packed arrays (large $N_{\rm ant}/\sqrt{u_{\rm max}}$), fftvis delivers runtime speedups of up to two orders of magnitude and significantly reduced memory usage compared to matvis.
@@
Throughout this work, we have identified several areas of future development that promise to further enhance the flexibility of fftvis. One straightforward addition to the simulator will be to extend fftvis to use GPU acceleration via cufinufft (Shih et al., 2021) and the existing GPU components that exist in the matvis codebase. This will potentially improve the performance of fftvis by an additional order-of-magnitude, making large-scale simulations even more tractable. Another potential area of improvement is removing the current limitation of supporting only a single, array-wide beam model by implementing per-antenna beam simulations. As discussed in Section 2.2.3, fftvis is expected to scale poorly with the number of unique beams requested by a user. However, this extension would offer the enhanced flexibility necessary for some systematic studies, and may remain more efficient than matvis if the number of unique beams is small. Finally, we also plan on exploring advanced techniques, such as the spectral basis decomposition outlined in Appendix A to reduce NUFFT calls for wide-band simulations of spectrally smooth sources and beams, which may also improve fftvis simulation efficiency.
@@ Acknowledgements
All timing benchmarks and simulated visibilities in this work were generated with the publicly-available fftvis package (v1.0.0). The fftvis source is hosted on GitHub at https://github.com/tyler-a-cox/fftvis. The analysis and simulation scripts (parameter files, run-scripts, etc.) used to produce the exact results in Figures 2–8 are available from the corresponding author upon reasonable request.
@@
The work in this paper was made possible by the following software packages: numpy (Harris et al., 2020), pyuvdata (Hazelton et al., 2017), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), finufft (Barnett et al., 2019), astropy (Astropy Collaboration et al., 2013, 2018), healpy (Zonca et al., 2019), and pygdsm (Price, 2016)
@@ Appendix A Efficient Evaluation of Spectrally Smooth Sky Models with fftvis
$\mathbf{v}_{k}=\sum_{i,j}\mathbf{s}_{i}\mathbf{b}_{j}\mathbf{d}_{ij,k},$ (15)
where $\mathbf{d}_{ij,k}$ are the structure constants of the basis $\mathbf{M}$. These constants reflect how the product of two basis functions can itself be expressed within the same basis, thus avoiding a computationally expensive reprojection. Once the product coefficients $\mathbf{v}_{k}$ are computed, a type-2 NUFFT (non-uniform to uniform) transforms this representation from a non-uniform to a uniform grid in the $uv$-plane. The resolution of this uniform grid is carefully chosen to satisfy the resolution constraints imposed by the highest observing frequency, defined as:

### v2

@@ 5.3 Memory Requirements
    NUFFT Intermediate Grid: The temporary uniform grid used internally by finufft. Its size in each dimension $i$, $n_{i}$, is determined by Equation 5, which depends on the maximum source extent (which we take to be $S_{i}=1$ for sky models that extend to the horizon), the maximum baseline coordinate, $X_{i}$, the desired precision, $\epsilon$, and oversampling factor, $\sigma$. The memory explicitly scales as $4\sigma N_{\rm grid}$.
@@
In matvis, a practical approach to managing memory involves chunking computations along the source axis. This technique allows for a nearly arbitrary reduction in the peak memory usage by processing subsets of sources sequentially, with generally minimal computational overhead. While the overhead might increase for very fine-grained chunking, it typically remains manageable for current hardware memory capacities. However, this source chunking strategy is generally less useful for fftvis when the maximum extent of the array is large. As shown in the memory tests above, the memory usage of fftvis is typically dominated by the size of the intermediate FFT grid. This grid size is not dependent on the number of sources, but rather on the maximum angular extent of the sky model being considered ($S_{i}$) and the maximum baseline coordinates ($X_{i}$). As a result, simply reducing the number of sources processed per chunk does not shrink the required grid size if the sources across all chunks collectively span the same wide angular extent. One potential way to mitigate poor memory scaling in fftvis would involve spatial chunking, or dividing the simulations into chunks based on the sources occupying localized regions of the sky. By limiting the angular extent ($S_{i}$) for sources within a single chunk, the factors determining the grid size in Equation 5 could be reduced, thus decreasing the maximum memory required for processing that chunk. Implementing such a scheme would require careful partitioning of the source catalog and potentially reinitializing the NUFFT for each spatial chunk, which may increase the computational runtime. We leave the development and evaluation of this spatial chunking method for future work.
@@
Figure 7: Comparison of the estimated memory usage of fftvis and matvis for the computational scaling cases presented in Figure˜5. For results shown here, we assume no chunking of the sky model in the processing stage, which can reduce the memory usage for matvis by summing over subsets of the sky model. We find that memory usage of fftvis is significantly less than that of matvis especially when the number of antennas and sources is large due to matvis forming an $N_{\rm sources}N_{\rm ants}$ matrix. When $u_{\max}$ is large, fftvis memory usage exceeds that of matvis due to the formation of the underlying fine FFT grid that scales with the physical size of the array in units of wavelengths. As with the computational scaling explored in Figure 5, the memory usage for arrays with a large $u_{\rm max}$ can be dramatically reduced if the array is gridded, as shown at in the right-hand column.
@@ 5.4 Multicore Performance and Parallelization
In fftvis, we adopt a hybrid strategy that prioritizes multiprocessing while still incorporating multithreading. Specifically, we parallelize the entire workflow across multiple processes, then assign the remaining cores as threads within each process. This approach is motivated by findings in 2025RASTI...4....1K, where a pure multithreaded strategy exhibited a performance plateau after just a few cores due to inherent scaling limitations and shared resource contention. In contrast, a multiprocessing strategy avoids many of these bottlenecks, providing more favorable scaling across a larger number of cores, albeit with a modest overhead associated with initializing processes.
To manage the distribution of these independent tasks across processes on a single machine, fftvis takes advantage of the ray library. Originally developed for scalable machine learning, ray provides a robust framework for efficiently scaling Python applications from a single core to many. Specifically, fftvis uses ray to divide the overall simulation workload by distributing the calculations across the $N_{\rm proc}$ available processes, assigning each process a roughly equal portion of the total workload determined by the number of simulation time steps and frequency channels. This approach helps manage the creation and coordination of multiple processes, simplifying the implementation of our chosen multiprocessing strategy. Furthermore, ray supports shared memory, enabling large, read-only objects like the beam, sky model, and model coordinates to be loaded once and accessed by multiple worker processes.
The total memory requirement of fftvis in this shared memory state can be approximated as
$M_{\mathrm{tot}}\approx M_{\mathrm{shared}}+N_{\mathrm{proc}}\,M_{\mathrm{local}},$ (17)
where $M_{\mathrm{shared}}$ represents arrays that can reside in shared memory (e.g., raw beam values and sky model components), and $M_{\mathrm{local}}$ denotes per-process allocations such as interpolated beam grids and per-source coordinate transformations. Explicitly,
$M_{\mathrm{shared}}$ $\simeq\underbrace{N_{\mathrm{sources}}\,N_{\nu}\,\beta_{\mathrm{f}}}_{\text{source flux}}+\underbrace{N_{\rm bls}\,N_{\nu}\,N_{\mathrm{feed}}^{2}\,N_{\rm times}\,\beta_{\mathrm{c}}}_{\text{visibility array}}$
$\ \ +\underbrace{N_{\nu}\,N_{\mathrm{beampix}}\,\beta_{\mathrm{c}}}_{\text{raw beam storage}},$ (18)
$M_{\mathrm{local}}$ $\simeq\underbrace{N^{2}_{\mathrm{feed}}\,N_{\mathrm{sources}}\,\beta_{\mathrm{c}}}_{\text{interpolated beam}}+\underbrace{3\,N_{\mathrm{sources}}\,\beta_{\mathrm{f}}}_{\text{source coordinates}}+\underbrace{N_{\mathrm{feed}}^{2}\,n_{x}\,n_{y}\,\beta_{\mathrm{c}}}_{\text{NUFFT grid}},$ (19)
where $\beta_{\mathrm{f}}$ and $\beta_{\mathrm{c}}$ are the byte sizes of real and complex double-precision values, respectively. In the current implementation, $M_{\mathrm{shared}}$ includes large arrays that are loaded once and accessed by all worker processes, whereas $M_{\mathrm{local}}$ scales linearly with $N_{\mathrm{proc}}$, increasing the total memory with the number of processes. While this shared memory capability is beneficial, a current limitation within fftvis’s implementation is that each process still independently interpolates the beam and performs coordinate rotations for all $N_{\rm sources}$ on each process independently. Therefore, these specific steps do not fully utilize memory sharing, requiring that arrays of size $\mathcal{O}\left(N_{\rm sources}\right)$ be stored in memory for each independent process.
@@
While the performance analysis presented here is restricted to a single computational node, the underlying fftvis framework is readily extensible to multi-node environments. In our current deployment on NMASCO and similar clusters, distributed execution is achieved by partitioning the total bandwidth into independent frequency blocks and submitting these as separate slurm array jobs, each running an instance of fftvis with shared configuration files. Because visibilities at different frequencies are computed independently, this strategy yields near-ideal scaling across nodes with a small amount of communication overhead. More generally, since fftvis is built on the ray distributed execution framework, native multi-node parallelism can be supported directly within the package. Future development will focus on integrating this capability into the standard workflow, allowing ray to manage task scheduling and inter-node communication transparently, streamlining large-scale simulations without reliance on external job arrays.
@@ 6 Conclusion
We have demonstrated through a series of validation tests that fftvis achieves numerical accuracy comparable to the established visibility simulator, matvis, itself validated against pyuvsim in 2025RASTI...4....1K. Our first set of comparisons against analytic solutions for diffuse sky models confirms the accuracy of the NUFFT implementation, while assessments of delay and fringe-rate spectra show that fftvis preserves the spectral and temporal coherence required for end-to-end validation tests. We find that the visibilities produced do not introduce significant spectral or temporal artifacts when operated at sufficient NUFFT precision ($\epsilon\lesssim 10^{-13}$). Furthermore, our benchmarking tests highlight the specific regimes where fftvis offers substantial advantages. For simulations involving large, densely-packed arrays (large $N_{\rm ant}/\sqrt{u_{\rm max}}$), fftvis delivers runtime speedups of up to two orders of magnitude and significantly reduced memory usage compared to matvis.
@@
Throughout this work, we have identified several areas of future development that promise to further enhance the flexibility of fftvis. One straightforward addition to the simulator will be to extend fftvis to use GPU acceleration via cufinufft (shih2021cufinufftloadbalancedgpulibrary) and the existing GPU components that exist in the matvis codebase. This will potentially improve the performance of fftvis by an additional order-of-magnitude, making large-scale simulations even more tractable. Another potential area of improvement is removing the current limitation of supporting only a single, array-wide beam model by implementing per-antenna beam simulations. As discussed in Section 2.2.3, fftvis is expected to scale poorly with the number of unique beams requested by a user. However, this extension would offer the enhanced flexibility necessary for some systematic studies, and may remain more efficient than matvis if the number of unique beams is small. Finally, we also plan on exploring advanced techniques, such as the spectral basis decomposition outlined in Appendix A to reduce NUFFT calls for wide-band simulations of spectrally smooth sources and beams, which may also improve fftvis simulation efficiency.
@@ Acknowledgements
All timing benchmarks and simulated visibilities in this work were generated with the publicly‑available fftvis package (v1.0.0). The fftvis source is hosted on GitHub at https://github.com/tyler-a-cox/fftvis. The analysis and simulation scripts (parameter files, run‑scripts, etc.) used to produce the exact results in Figures 2–8 are available from the corresponding author upon reasonable request.
Conflict of Interest
Authors declare no conflict of interest.
Code Availability
@@
The work in this paper was made possible by the following software packages: numpy (harris2020array), pyuvdata (2017JOSS....2..140H), scipy (2020SciPy-NMeth), matplotlib (Hunter:2007), finufft (barnett2019parallelnonuniformfastfourier), astropy (2013A&A...558A..33A; 2018AJ....156..123A), healpy (2019JOSS....4.1298Z), and pygdsm (2016ascl.soft03013P)
@@ Appendix A Efficient Evaluation of Spectrally Smooth Sky Models with fftvis
$\mathbf{v}_{k}=\sum_{i,j}\mathbf{s}_{i}\mathbf{b}_{j}\mathbf{d}^{k}_{ij},$ (23)
where $\mathbf{d}^{k}_{ij}$ are the structure constants of the basis $\mathbf{M}$, defined as
$\mathbf{d}_{ij}^{k}=\sum_{q}\mathbf{M}_{qi}\mathbf{M}_{qj}\mathbf{M}_{qk}.$ (24)
These constants reflect how the product of two basis functions can itself be expressed within the same basis, thus avoiding a computationally expensive reprojection. Once the product coefficients $\mathbf{v}_{k}$ are computed, a type-2 NUFFT (non-uniform to uniform) transforms this representation from a non-uniform to a uniform grid in the $uv$-plane. The resolution of this uniform grid is carefully chosen to satisfy the resolution constraints imposed by the highest observing frequency, defined as: